# PinnDE: Physics-Informed Neural Networks for Solving Differential Equations

### v1

@@ 4.1 System of ordinary differential equations
4.2 Linear advection equation
We next use PinnDE to solve the linear advection equation with a standard PINN which uses a soft constrained initial condition. We use periodic boundary conditions, which in PinnDE are implemented at the neural network level to imposes periodic boundaries on the independent spatial variable $x$. This removes the boundary component of the composite loss function described in Eqn. (2) when using periodic boundaries. The linear advection with advection speed $c=1$ is described as
$\frac{\partial u}{\partial t}+\frac{\partial u}{\partial x}=0,$ (8a)
and we use the initial condition
$u_{0}(x)=\cos\pi x.$ (8b)
We solve this equation over the interval $t\in[0,1]$, on the spatial domain $x\in[-1,1]$. We compare the neural network solution to the analytical solution $u(t,x)=\cos\pi(x-t)$ in Figure 4, showing the success of the trained model, as it gives small point-wise and overall mean-squared errors, respectively.
Figure 4: Solution predicted using a PINN for the linear advection equation (8) (left), the analytical solution (middle), and the mean squared error between them (right).
For this example we used a PINN with 4 hidden layers with 60 nodes per layer, the hyperbolic tangent activation function, and trained the neural network using the Adam optimizer. A total of 100 initial value collocation points were used for the network to learn the initial condition, and 10000 collocation points were used across the spatio-temporal domain for the network to learn the solution, both sampled using Latin hypercube sampling. The minimum of the loss was reached after 5000 epochs of training.
The code for solving this equation can be found in Appendix A.2.
4.3 Poisson equation
We next use PinnDE to solve a specific instance of the Poisson equation with a PINN which uses hard constrained boundary conditions. This forces boundary values to be satisfied by the neural network exactly, leading to zero errors at the boundaries. We solve this equation using Dirichlet boundary values. The particular form of this equation we solve is
$\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}=-2\pi^{2}\cos\pi x\sin\pi y,$ (9)
with the Dirichlet boundary conditions being obtained from the exact solution for this problem given by $u(x,y)=\cos\pi x\sin\pi y.$ We solve this equation over the square domain $(x,y)\in[-1,1]\times[-1,1]$. We compare the neural network solution to the analytical solution in Figure 5.
Figure 5: Solution predicted using a PINN for the Poisson equation (9)
(left), the analytical solution (middle), and the mean squared error between them (right).
This solution was obtained by training a PINN with 5 hidden layers with 40 nodes per layer, using the hyperbolic tangent as activation function and Adam as the chosen optimizer. A total of 10000 collocation points were used across the domain for the network to learn the solution, sampled using Latin hypercube sampling. As the network was hard constrained, no points along the boundaries needed to be used to learn the boundary values, as these are enforced automatically. The loss reached a minimum after 5000 epochs by which point we stopped training.
@@ 4.4 Heat Equation
We lastly use PinnDE to solve the linear heat equation with a deep operator network which uses soft constrained initial and boundary conditions, meaning we use a composite loss function consisting of the PDE, initial condition, and boundary condition losses. We solve this equation with Dirichlet boundary values. Specifically, we consider the equation
$\frac{\partial u}{\partial t}=\nu\frac{\partial^{2}u}{\partial x^{2}},\quad\nu=0.1,$ (10a)
@@
$u(t,0)=u(t,1)=0.$ (10c)
We solve this equation over the interval $t\in[0,1]$, on the spatial domain $x\in[0,1]$. We compare the neural network solution to the analytical solution $u(t,x)=\exp(\nu\pi^{2}t)\sin\pi x$ in Figure 6.
Figure 6: Solution predicted by DeepONet for the heat equation (10)
(left), the analytical solution (middle), and the mean squared error between them (right).
For this example we trained a physics-informed DeepONet with 4 hidden layers with 60 nodes per layer, using the hyperbolic tangent activation function and Adam as our optimizer. 100 initial value points were used to be sampled along $t=0$ for the network to learn the initial condition. The soft-constrained boundary was sampled using 100 boundary value points along $x=0$ and $x=1$ for the neural network to learn the boundary conditions, and 10000 collocation points were used across the domain for the network to learn the solution itself, both sampled with Latin hypercube sampling. We sample 30000 initial conditions for the DeepONet within a range of $u_{0}(x)\in[-2,2$]. We train this network for 3000 epochs which is where we have already reached a minimum loss.
@@ 5 Conclusion
In this paper we presented PinnDE, an open-source software package in Python for PINN and DeepONet implementations for solving ordinary and partial differential equations. We reviewed the methodologies behind PINN and DeepONet architectures, then gave a summary of PinnDE and its functionality, then presented several worked examples to show the effectiveness of this package. In PinnDE, we highlight the overall structure of the package and display the methods available for a user to solve differential equations using physics-informed neural networks. In contrast to other packages, PinnDE stresses simplicity and requires minimal coding from the side of the user, thus alleviating the need of writing higher amounts of software to support the higher level of customizability other packages offer. As such, PinnDE provides a simple interface where users with little experience can still write and understand the implementation of software written in this package. This gives researchers the ability to use this package with collaborators outside of the field who may be interested in this area of research but lacking the skills to use a more advanced package. This also provides a tool for educators to possibly use as the simple implementations gives a low barrier of entry to new users.
PinnDE is continuously being developed to integrate new advancements in research surrounding PINNs, while aiming to maintain a straightforward user experience. Some possible additions which will be implemented in the future are adaptive collocation point methods, other variants on DeepONet and PINN architectures, meta-learned optimization, the availability of different backends, and general geometries which differential equations can be solved over.
@@ Appendix A Appendix
Here we provide the code corresponding to the examples presented in Sections [4.1, 4.2, 4.3, 4.4].
A.1 System of ordinary differential equations code
@@
    import pinnde.ode_Solvers as ode_Solvers
@@
    eqns = ["utt + u", "xt+u"]
    orders = [2, 1]
    inits = [[0.5, 1],[2]]
@@
    N_sensors = 5000
    epochs = 20000
    mymodel = ode_Solvers.solveODE_DeepONetSystem_IVP(eqns, orders, inits,
                            t_bdry, N_pde, sensor_range, N_sensors, epochs)
    exact_eqns = ["np.sin(t)+0.5*np.cos(t)", "-0.5*np.sin(t)+np.cos(t)+1"]
    mymodel.plot_predicted_exact(exact_eqns)
@@
    mymodel.timeStep(10)
A.2 Linear advection equation code
@@
    import pinnde.pde_Solvers as pde_Solvers
    import pinnde.pde_Initials as pde_Initials
    import pinnde.pde_Boundaries_2var as pde_Boundaries
@@
    u0 = lambda x: tf.cos(np.pi*x)
    t_bdry = [0,1]
    x_bdry = [-1,1]
    t_order = 1
    N_iv = 100
    initials = pde_Initials.setup_initials_2var(t_bdry, x_bdry, t_order,
                                                              [u0], N_iv)
    boundaries=pde_Boundaries.setup_boundaries_periodic_tx(t_bdry,
                                                            x_bdry)
    eqn = "ut+ux"
    N_pde = 10000
    epochs = 5000
    mymodel = pde_Solvers.solvePDE_tx(eqn, initials,
                                    boundaries, N_pde)
    mymodel.train_model(epochs)
    mymodel.plot_predicted_exact("tf.cos(np.pi*(x-t))")
A.3 Poisson equation code
@@
    import pinnde.pde_Solvers as pde_Solvers
    import pinnde.pde_Boundaries_2var as pde_Boundaries
@@
    x_bdry = [-1, 1]
    y_bdry = [-1, 1]
    N_bc = 100
    all_boundary = lambda x, y: tf.cos(np.pi*x)*tf.sin(np.pi*y)
    boundaries = pde_Boundaries.setup_boundaries_dirichlet_xy(x_bdry,
                    y_bdry,  N_bc, all_boundaries_cond=all_boundary)
    eqn = "uxx + uyy - (-2*np.pi**2*tf.cos(np.pi*x)*tf.sin(np.pi*y))"
    N_pde = 10000
    epochs = 5000
    mymodel = pde_Solvers.solvePDE_xy(eqn, boundaries, N_pde,
                    net_layers=5, net_units=40, constraint = "hard")
    mymodel.train_model(epochs)
    mymodel.plot_predicted_exact("tf.cos(np.pi*x)*tf.sin(np.pi*y)")
A.4 Heat equation code
@@
    import pinnde.pde_Solvers as pde_Solvers
    import pinnde.pde_Initials as pde_Initials
    import pinnde.pde_Boundaries_2var as pde_Boundaries
@@
    u0 = lambda x: tf.sin(np.pi*x)
    t_bdry = [0,1]
    x_bdry = [0,1]
    t_order = 1
    N_iv = 100
    initials = pde_Initials.setup_initials_2var(t_bdry, x_bdry, t_order,
                                                              [u0], N_iv)
    all_boundary = lambda x: 0+x*0
    N_bc = 100
    boundaries = pde_Boundaries.setup_boundaries_dirichlet_tx(t_bdry,
                        x_bdry, N_bc, all_boundaries_cond=all_boundary)
    eqn = "0.1*uxx - ut"
    N_pde = 10000
    N_sensors = 30000
    sensor_range = [-2, 2]
    epochs = 3000
    mymodel = pde_Solvers.solvePDE_DeepONet_tx(eqn, initials, boundaries,
                                    N_pde, N_sensors, sensor_range)
    mymodel.train_model(epochs)
    exact_eqn = "np.e**(-((np.pi**2) * 0.1 * t))*np.sin(np.pi*x)"
    mymodel.plot_predicted_exact(exact_eqn)

### v2

@@ 4.3 Dam break problem
4.4 Burgers’ equation
We next use PinnDE to solve Burgers’ equation with a PINN which uses soft-constrained initial and boundary conditions. We solve this equation using periodic boundary values. We also use a residual-based adaptive distribution (RAD) adaptive sampling strategy. In particular, we consider
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}-\nu\frac{\partial^{2}u}{\partial x^{2}}=0,\qquad\nu=\frac{0.01}{\pi}.$ (13)
We solve this equation over the square domain $(t,x)\in[0,1]\times[-1,1]$. We compare the neural network solution to a numerical reference solution in Figure 6. The numerical reference solution uses a spectral discretization in space combined with the method of lines based on the Runge–Kutta integrator for time-stepping.
Figure 6: Solution predicted using a PINN for Burgers’ equation (13) (left), the numerical reference solution (middle), and the mean squared error between them (right).
For this example we used a PINN with 4 hidden layers with 60 nodes per layer, the hyperbolic tangent activation function, and trained the neural network using a two step training routine. The Adam optimizer was used to train the network for 5000 epochs. A total of 400 initial value collocation points were used for the network to learn the initial condition, and 2000 collocation points were used across the spatial domain for the network to learn the solution, all sampled using Latin hypercube sampling.
We used residual-based adaptive distribution (RAD) to train this equation. This strategy resamples the distribution of collocation points during training to align with areas of larger PDE residual. For Burgers’ equation, this is particularly effective as we have a steep developing shock which to resolve requires locally more collocation points. We resample the distribution after 3000 epochs of training. We sample 10000 collocation points across the domain, and sample 2000 new collocation points from the probability distribution function described in Section 2.4, with hyperparameters $k=3,c=0$. We compare these points in Fig. 7.
Figure 7: Original collocation points sampled using Latin hypercube sampling (right) and the resampled distribution by RAD (left).
@@ 4.5 Heat Equation
For our next example we use PinnDE to solve the $(1+2)$-dimensional linear heat equation with a DeepONet which uses soft-constrained initial and boundary conditions, meaning we use a composite loss function consisting of the PDE, initial condition, and boundary condition losses. We solve this equation with Dirichlet boundary values. Specifically, we consider the equation
$\frac{\partial u}{\partial t}=\nu\left(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\right),\quad\nu=0.15,$ (14a)
@@
$u(t,x,y)=0\quad\forall(t,x,y)\in\{[0,1]\times\mathbb{R}^{2}|x^{2}+y^{2}=1\}$ (14c)
We solve this equation over the time interval $t\in[0,1]$, on the spatial domain of the unit circle $(x,y)\in\{\mathbb{R}^{2}|x^{2}+y^{2}<1\}$. We compare the neural network solution to the analytical solution $u(t,x,y)=\exp(\nu\pi^{2}t)\sin\pi x\sin\pi y$ in Figure 8.
Figure 8: Solution predicted by DeepONet for the heat equation (14) (left), the analytical solution (middle), and the mean squared error between them (right).
For this example we trained a DeepONet with both branch net and trunk net being multi-layer perceptrons having 4 hidden layers with 60 nodes per layer, using the hyperbolic tangent activation function. We use Adam as our chosen optimizer. 600 initial value points were used to be sampled along $t=0$ for the network to learn the initial condition. The soft-constrained boundary was sampled using 600 boundary value points for the neural network to learn the boundary conditions, and 12000 collocation points were used across the domain for the network to learn the solution itself, both sampled with Latin hypercube sampling. We sample a total of 1000 initial conditions from a class of truncated Fourier series with random coefficients to learn the solution operator. After training on this class of initial conditions, the operator network can be evaluated in a simple inference step for new initial conditions from the same initial condition class, an example of which is shown in Figure 8. We note here that different classes of initial conditions, such as Gaussian random fields lu2019deeponet could be considered as well for operator learning, and will be implemented in PinnDE in the future. We train this network for 4000 epochs which is when the minimum loss is reached
The code for solving this equation can be found in Appendix A.5.
4.6 Inverse linear advection speed
We lastly use PinnDE to solve for an unknown advection speed in the linear advection equation with an inverse PINN which uses a soft-constrained initial condition. We use periodic boundary conditions, which in PinnDE are implemented at the neural network level as a hard constraint to imposes periodic boundaries on the independent spatial variable $x$. We also use a residual-based adaptive refinement with distribution (RAR-D) adaptive sampling strategy. The linear advection with variable advection speed for our reference solution data is $c=\sin(\pi x)$, and the equation is described as
$\frac{\partial u}{\partial t}+c\frac{\partial u}{\partial x}=0,$ (15a)
where we use the initial condition
$u_{0}(x)=\cos\pi x.$ (15b)
We solve this equation over the interval $t\in[0,1]$, on the spatial domain $x\in[-1,1]$. We compare the neural network solution to a numerical calculation of the solution in Figure 9, showing the success of the trained model, as it gives small point-wise and overall mean-squared errors, respectively. We use a numerical solution here as no analytical solution for this equation exists.
Figure 9: Solution predictions using a PINN for the inverse linear advection speed problem (15) (left), the numerical solution (middle), and the difference between them (right). Top: $u(t,x)$. Bottom: Advection velocity $c=\sin(\pi x)$.
For this example we used a PINN with 4 hidden layers with 60 nodes per layer, the hyperbolic tangent activation function, and trained the neural network using the Adam optimizer. A total of 200 initial value collocation points were used for the network to learn the initial condition, and 1000 collocation points were used across the spatio-temporal domain for the network to learn the solution, both sampled using Latin hypercube sampling. We also sampled 500 points across the spatio-temporal domain over which we evaluate the analytical solution $u$ as data points. These data points are then used to learn the advection speed $c$. Our RAR-D sampling strategy added 30 collocation points every 250 epochs, with hyper-parameters $k=3,c=0$ (see wu2023comprehensive for more details). The evolution of collocation points throughout training is shown in Fig. 10. The minimum of the loss was reached after 5000 epochs of training.
Figure 10: Original collocation points before training (left), the collocation points halfway through training at epoch 2500 (middle), and the final set of collocation points after 5000 epochs (right).
@@ 5 Conclusion
Deep learning-based approaches for solving differential equations have become an increasingly popular alternative for solving differential equations compared to methods based around numerical differentiation such as finite element methods, finite volume methods, and finite difference methods. PINNs and DeepONets are two methodologies of this deep learning-based approach. With these, we have a naturally meshless alternative to any of the classical numerical methods mentioned above. Deep learning-based approaches have the ability to use automatic differentiation to compute derivatives, without the need for any numerical discretizations to approximate derivatives. This streamlines the application of these methods across arbitrary domains, and by being able to incorporate physical constraints into the solving procedure, allow for them to solve problems in which classic numerical methods may struggle.
In this paper we presented PinnDE, an open-source software package in Python for PINN and DeepONet implementations for solving ordinary and partial differential equations. We reviewed the methodologies behind physics-informed neural networks for both forward and inverse problems, the DeepONet approach to operator learning, and adaptive point sampling strategies. We then introduced PinnDE and its functionality, and finally presented several worked examples to show the effectiveness of this package.
Software packages giving the ability to use physics-informed machine learning approaches are becoming increasingly popular. In contrast to other existing packages, PinnDE stresses simplicity and requires minimal coding from the side of the user, thus alleviating the need of writing significant amounts of software to support the higher level of customization that other packages offer, while still providing many of the abilities that alternative packages offer. As such, PinnDE provides a simple interface where users with little experience can still write and understand the implementation of software written in this package. This gives researchers the ability to use this package with collaborators outside of the field who may be interested in this area of research but are lacking the skills to use more advanced packages. PinnDE also provides a tool for educators to use as the simple implementations give a low barrier of entry to new users. We also hope that by providing a package with such quick implementations of problems will promote further research and development of the field of scientific machine learning.
PinnDE is continuously being developed to integrate new advancements in research surrounding PINNs and DeepONets, while aiming to maintain a straightforward user experience. Some possible additions which will be implemented in future versions are more adaptive collocation point methods, other variants on DeepONet and PINN architectures, meta-learned optimization, the availability of different backend, and general geometries which differential equations can be solved over.
@@ Appendix A Appendix
Here we provide the code corresponding to the examples presented in Sections [4.1, 4.2, 4.3, 4.4, 4.5, 4.6].
A.1 Rigid body code
@@
    import src.pinnde as p
@@
    re = p.domain.NRect(1, [0], [2.5])
    cond = p.boundaries.odeicbc(tre, [[1], [1], [1]], "ic")
    dat = p.data.pinndata(tre, cond, 3000, 300)
    eqns = ["(-5/6)*u2*u3 - u1x1", "(5/2)*u1*u3 - u2x1",
                                    "(-5/3)*u1*u2 -u3x1"]
    mymodel = p.models.pinn(dat, eqns)
    mymodel.train(20000)
    mymodel.train(5000, opt="lbfgs")
A.2 Lorenz model code
    import pinnde.legacy.ode_Solvers as ode_Solvers
    eqn1 = "(-1.6)*x*y - ut"
    eqn2 = "0.1*u*y - xt"
    eqn3 = "0.75*u*x - yt"
    eqns = [eqn1, eqn2, eqn3]
    orders = [1, 1, 1]
    inits = [[0.5], [0.75], [1]]
@@ A.2 Lorenz model code
    num_sensors = 3000
    epochs = 75000
    mymodel = ode_Solvers.solveODE_DeepONetSystem_IVP(eqns, orders,
            inits, t_bdry, N_pde, sensor_range, num_sensors, epochs)
    mymodel.timeStep(6)
@@
    mymodel.plot_solution_prediction()
A.3 Dam break problem code
@@
    import pinnde as p
@@
    h_analytical = lambda t, x: tf.where(x+0.25<-0.5*t, 0.5**2,
     tf.where(x+0.25<1*t, (1-(x+0.25)/t)**2/(9), 0.0))
    u_analytical = lambda t, x: tf.where(x+0.25<-0.5*t, 0.0,
        tf.where(x+0.25<1*t, 2/3*(0.5+(x+0.25)/t), 0.0))
    h_init = lambda x: tf.where(x+0.25<-0.5*0.75, 0.5**2,
        tf.where(x+0.25<1*0.75, (1-(x+0.25)/0.75)**2/(9), 0.0))
    u_init = lambda x: tf.where(x+0.25<-0.5*0.75, 0.0,
        tf.where(x+0.25<1*0.75, 2/3*(0.5+(x+0.25)/0.75), 0.0))
    tre = p.domain.Time_NRect(1, [-np.pi], [np.pi], [0.01, 1])
    bound = p.boundaries.dirichlet(tre, [h_analytical])
    inits = p.initials.initials(tre, [[h_init], [u_init]])
    dat = p.data.timepinndata(tre, bound, inits, 15000, 2000, 1000)
    eqn1 = "u1*u2t + u1t*u2 + 2*u1*u2*u2x1 + u1x1*u2**2 + u1*u1x1"
    eqn2 = "u1t + u1*u2x1 + u1x1*u2"
    mymodel = p.models.pinn(dat, [eqn1, eqn2])
    mymodel.train(35000)
A.4 Burgers’ equation code
@@
    import pinnde as p
@@
    tre = p.domain.Time_NRect(1, [-1], [1], [0,1])
    bound = p.boundaries.periodic(tre)
    inits = p.initials.initials(tre, [lambda x: -np.sin(np.pi*x)])
    dat = p.data.timepinndata(tre, bound, inits, 2000, 600, 400)
    mymodel = p.models.pinn(dat, ["ut+u*ux1-(0.01/np.pi)*ux1x1"])
    rad = p.adaptives.RAD(3000, 3, 0, 5)
    mymodel.train(5000)
A.5 Heat equation code
@@
    import pinnde as p
@@
    center = [0.5, 0.5]
    semilengths = [0.5, 0.5]
    timerange = [0, 1]
    dom = p.domain.Time_NEllipsoid(2, center, semilengths, timerange)
    bdryfunc = lambda t, x1, x2: 0+t*0
    bound = p.boundaries.dirichlet(tre, [bdryfunc])
    u0func = lambda x1, x2: tf.sin(np.pi*x1)*tf.sin(np.pi*x2)
    inits = p.initials.initials(dom, [u0func])
    dat = p.data.timedondata(dom, bound, inits, 12000, 600, 600, 1000)
    eqn = "0.15*ux1x1 + 0.15*ux2x2 - ut"
    mymodel = p.models.deeponet(dat, [eqn])
    mymodel.train(4000)
A.6 Inverse linear advection equation code
    import src.pinnde as p
    import numpy as np
    import tensorflow as tf
    from scipy.integrate import odeint
    from scipy.fftpack import diff as psdiff
    from scipy.interpolate import RegularGridInterpolator
    c = lambda x: tf.sin(np.pi*x)
    u0 = lambda x: tf.cos(np.pi*x)
    # Creating reference solution
    t = np.linspace(0, 1, 200)
    x = np.linspace(-1, 1, 200)
    c_val = c(x)
    def la(u, t, L):
      ux = psdiff(u, period=L)
      ut = -c_val*ux
      return ut
    def solve_la(u0, t, L):
      sol = odeint(la, u0, t, args=(L,))
      return sol
    ureference = solve_la(u0(x), t, 2)
    uinterp = RegularGridInterpolator((t, x), ureference,
                method=’linear’)
    N_data = 500
    xdata = np.random.uniform(-1, 1, N_data)
    tdata = np.random.uniform(0, 1, N_data)
    data = np.column_stack([tdata, xdata])
    udata = uinterp(data)
    timerange = [0, 1]
    tre = p.domain.Time_NRect(1, [-1], [1], timerange)
    bound = p.boundaries.periodic(tre)
    u0func = lambda x1: tf.cos(np.pi*x1)
    inits = p.initials.initials(tre, [u0])
    dat = p.data.timeinvpinndata(tre, bound, inits, [tdata, xdata],
                                [udata], 1000, 10, 200)
    eqn = "ut+c*ux1"
    mymodel = p.models.invpinn(dat, [eqn], ["c"])
    rard = p.adaptives.RARD(250, 30, k=3, c=0)
    mymodel.train(5000, adapt_pt=rard)